OpenStudy (anonymous):

How do I find length, width, and height of a rectangle if all I am given is the volume?

OpenStudy (anonymous):

i don't know

OpenStudy (anonymous):

nope. Just the volume given is 216

OpenStudy (anonymous):

thanks anyway.

OpenStudy (anonymous):

yw

jhonyy9 (jhonyy9):

for example you can calculi adding values or for example 216/4 =54 what mean that the height is equal 4 so than area of base is equal 54 what you can getting if the rectangle has a width of 6 and a length of 9 because 9*6=54*4=216